A 3 by 3 by 3 cube has three holes, each with a 1 by 1 cross-section running from the centre of
each face to the centre of the opposite face. What is the total surface area (in square units) of the
resulting solid?

Answer by ikleyn(52781)

.

The surface area of the original cube is/was  6*(3*3) = 6*9 = 54 square unit.


After 3 holes were made, we should subtract  6*(1*1) = square units for/(or from) 6 faces;

and add 6 times 4 square units (in total 6*4 = 24 square units) for six  1 x 1 x 1  small empty cubes.


So, the ANSWER is  54 - 6 + 24 = 72 square units.